OpenStudy (anonymous):

integral calculus: y = the integral from -2 to root x of (cos(t)/ t^11) dt Then dy/dt = ?

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

@calculusxy try this one

OpenStudy (calculusxy):

Woah... So hard..

OpenStudy (anonymous):

which level are you in? College?

OpenStudy (calculusxy):

Middle school

OpenStudy (anonymous):

Nice man I am in college

OpenStudy (calculusxy):

Great.. Can you help me with my question?

OpenStudy (anonymous):

sure tag me in

OpenStudy (dumbcow):

i think its \[\frac{dy}{dt} = f(\sqrt{x})\frac{dx}{dt}\]

OpenStudy (anonymous):

and what does dx/dt equal to in this case? this is what I am looking for

OpenStudy (dumbcow):

hmm not sure we have enough info

OpenStudy (anonymous):

I checked with webwork the answer is incorret. The first part says that we should use part 1 of the fundamental theorem of calculus.

OpenStudy (dumbcow):

since y is given as function of x , to know how y is changing wrt "t" you need to know how x is changing wrt "t"

OpenStudy (anonymous):

What I wrote is the only information given. This is what has been dazzling me all day.

hartnn (hartnn):

\(\huge \dfrac{d}{dx}\int \limits_a^x f(t)dt = f(x) \\ \huge \dfrac{d}{dx}\int \limits_a^{g(x)}f(t)dt =f(g(x))\dfrac{d}{dx}g(x) \) i think you need dy/dx and not dy/dt ? because y will be the function of x and not t

OpenStudy (anonymous):

oh yes you are right I need dy/dx. I am sorry about that

OpenStudy (dumbcow):

haha

hartnn (hartnn):

good, so it'll just be f(sqrt x) d/dx sqrt x with f(t) = cos t/ t^11

OpenStudy (anonymous):

on webwork it says that dx has no contect this means that its not part of the solution.

hartnn (hartnn):

dx ? could you find f(sqrt x) ? do you how to get d/dx sqrt x ?

hartnn (hartnn):

if f(t) = cos t / t^11 whats f(sqrt x) ? can u find ?

OpenStudy (anonymous):

hey man I entered this into webwork and it says that the answer is equivalent to one that I submitted cos (sqrt(x) ) / (sqrt(x)^11)

hartnn (hartnn):

but i said, its this : "f(sqrt x) d/dx sqrt x" you only input f(sqrt x) part